Determine whether each relation defines a function
x-4y=8
y=2x+3
1/2x=3

Answer by ikleyn(52797)

.
Determine whether each relation defines a function
(a) x-4y=8
(b) y=2x+3
(c) 1/2x=3
~~~~~~~~~~~~~~~~

(a)  Yes, because you can express y as a function of x.


(b)  Yes, because you can express y as a function of x.


(c)  No, since it is not a  " relation "  in a mathematical sense,
     and it does not define any y-values.

Solved, answered and explained.